# Cascaded Multi-Scale Attention for Enhanced Multi-Scale Feature Extraction and Interaction with Low-Resolution Images

### v1

@@ 4.3.1 Training
Table 4: Results on CIFAR-100 with input size $32\times 32$. means the values reported in the respective papers.
Model Params CIFAR-100
(M) (%)
DeiT-B  (Touvron et al., 2021) 85.1 70.5
T2T-ViTt-14  (Yuan et al., 2021a) 21.1 75.9
Swin-B  (Liu et al., 2021) 86.7 78.5
PVTv2-B1  (Wang et al., 2022b) 13.5 80.5
P2T-L (Wu et al., 2022) 53.9 81.8
ConvNeXt-B  (Liu et al., 2022) 87.7 82.0
ResNet-50 (He et al., 2016) 23.5 82.9
SG-Former-M (Ren et al., 2023) 33.5 83.1
SMT-B (Lin et al., 2023b) 31.6 83.5
Shunted-B (Ren et al., 2022) 39.1 83.5
SG-Former-B (Ren et al., 2023) 77.1 83.8
SBCFormer-L  (Lu et al., 2024) 18.2 84.1
CMSA-S 4.2 84.5
CMSA-B 5.7 85.0
CMSA-L 7.4 85.2
4.3.2 Results
@@
Figure 4 and Table 4 present the performance of our models across varying parameter sizes compared to representative methods in CNNs, ViTs, and their hybrids. It shows averaged values over 300 trials. Notably, our largest model, CMSA-L, achieves 98.0% accuracy on CIFAR-10 and 85.2% accuracy on CIFAR-100, setting new benchmarks when limited to learning less than 300 epochs.
4.4 Ablation Study
@@
Table 5: Ablation tests of key components of CMSA for human pose estimation on COCO 2017 with $32\times 24$ inputs.
@@
Table 5 presents the results. The model in the first row, which has a hierarchical pyramid structure identical to our proposed network as shown in Figure 2a but uses standard self-attention instead of CMSA, performs the worst. The second-row model, while maintaining the same architecture, incorporates grouped multi-scale attention but lacks cascaded feature propagation, resulting in moderate performance improvements over the first model. The third-row model includes cascaded feature propagation but lacks other components, specifically removing CF and SF from the full model. The fourth-row model includes CF, and the fifth-row model includes both CF and SF, completing the network with CMSA. The performance improvements in these models demonstrate the effectiveness of the added components.
5 Conclusion
In this paper, we address the challenge of making visual inference on target objects from their low-resolution images, a critical issue in scenarios where targets are captured at low resolutions, such as with distant surveillance cameras or on low-performance edge devices. To mitigate the limitations inherent in low-resolution inputs—specifically, the difficulty in managing multi-scale features without compromising valuable information—we developed a novel attention mechanism, termed cascaded multi-scale attention (CMSA). This mechanism leverages a combination of multi-scale feature extraction and feature interactions, sidestepping the conventional downsampling operations for obtaining multi-scale features that typically result in loss of information in low-resolution contexts.
The empirical validation of our method across three tasks, i.e., human pose estimation, head pose estimation, and image classification, showed that our method surpasses current state-of-the-art approaches tailored to these applications, achieving this with fewer parameters. These results underscore the potential of our proposed method to significantly improve image recognition tasks under constrained conditions, marking a step forward in the deployment of deep learning models on edge devices and in situations where only low-resolution images are available. In conclusion, our research contributes a useful framework for enhancing the accuracy of image recognition in less-than-ideal conditions where high-resolution data is not accessible.
Acknowledgments

### v3

@@ IV-C1 Training
TABLE IV: Results on CIFAR-100 with input size $32\times 32$. means the values reported in the respective papers.
Model Params FLOPs CIFAR-100
(M) (G) (%)
DeiT-B [49] 85.1 21.8 70.5
T2T-ViTt-14 [67] 21.1 1.4 75.9
Swin-B [7] 86.7 5.0 78.5
P2T-B [13] 35.6 2.1 79.6
ConvNeXt-T [68] 27.8 1.5 80.3
PVTv2-B1 [38] 13.5 0.7 80.5
P2T-L [13] 53.9 3.5 81.8
ConvNeXt-B [68] 87.7 5.0 82.0
ResNet-50 [69] 23.5 1.3 82.9
SG-Former-M [8] 38.1 2.2 83.1
SMT-B [16] 31.6 2.5 83.5
Shunted-B [14] 39.1 2.2 83.5
SG-Former-B [8] 77.1 4.7 83.8
SBCFormer-L [17] 18.2 3.2 84.1
CMSA-S 4.2 0.9 84.5
CMSA-B 5.7 1.2 85.0
CMSA-L 7.4 1.5 85.2
IV-C2 Results
@@
Figure 4 and Table IV present the performance of our models across varying parameter sizes compared to representative methods in CNNs, ViTs, and their hybrids. It shows averaged values over 300 trials. Notably, our largest model, CMSA-L, achieves 98.0% accuracy on CIFAR-10 and 85.2% accuracy on CIFAR-100, setting new benchmarks when limited to learning less than 300 epochs.
Figure 5: Heatmap visualizations of the left wrist, elbow, and knee from models trained for human pose estimation with $32\times 24$ input. Redder regions indicate higher attention for the corresponding keypoint type in the heatmap.
Figure 6: Attention map visualizations from CMSA modules at each stage of CMSA-L for human pose estimation with input size $32\times 24$. Red points in inputs denote the corresponding 17 keypoints.
IV-D Visualization and Analysis
To further validate the effectiveness of the proposed CMSA, we visualize the predicted keypoint heatmaps generated by CMSA-L and methods using typical backbones with multi-scale representations for human pose estimation on the COCO 2017 dataset. As illustrated in Fig. 5, our model produces more accurate and sharply focused activations around body keypoints compared to existing multi-scale attention methods. Furthermore, we visualize the attention maps from the final CMSA module of each stage in CMSA-L in Fig. 6. The results indicate that, in the early stages, each keypoint attends broadly to other keypoints to capture a comprehensive global context. As the network deepens, the attention progressively focuses on smaller relevant regions, enhancing heatmap precision and keypoint localization. This contributes to the overall effectiveness of CMSA, particularly for low-resolution inputs.
IV-E Ablation Study
@@
Table V presents the results. The model in the first row, which has a hierarchical pyramid structure identical to our proposed network as shown in Fig. 2a but uses standard self-attention instead of CMSA, performs the worst. The second-row model, while maintaining the same architecture, incorporates grouped multi-scale attention but lacks cascaded feature propagation, resulting in moderate performance improvements over the first model. The third-row model includes cascaded feature propagation but lacks other components, specifically removing CF and SF from the full model. The fourth-row model includes SF, and the fifth-row model includes both CF and SF, completing the network with CMSA. The performance improvements in these models demonstrate the effectiveness of the added components.
TABLE V: Ablation tests of key components of CMSA for human pose estimation on COCO 2017 with $32\times 24$ inputs.
@@
V Conclusion
In this paper, we address the challenge of making visual inference on target objects from their low-resolution images, a critical issue in scenarios where targets are captured at low resolutions, such as with distant surveillance cameras or on low-performance edge devices. To mitigate the limitations inherent in low-resolution inputs—specifically, the difficulty in managing multi-scale features without compromising valuable information—we developed a novel attention mechanism, termed cascaded multi-scale attention (CMSA). This mechanism leverages a combination of multi-scale feature extraction and feature interactions, sidestepping the conventional downsampling operations for obtaining multi-scale features that typically result in loss of information in low-resolution contexts. The empirical validation of our method across three tasks, i.e., human pose estimation, head pose estimation, and image classification, showed that our method surpasses current state-of-the-art approaches tailored to these applications, achieving this with fewer parameters. These results underscore the potential of our proposed method to significantly improve image recognition tasks under constrained conditions, marking a step forward in the deployment of deep learning models on edge devices and in situations where only low-resolution images are available. In conclusion, our research contributes a useful framework for enhancing the accuracy of image recognition in less-than-ideal conditions where high-resolution data is not accessible.